of equation Solve the system 3x + 4y = -14 5x+7 y = -25
Answer #1

3x +4y=-14 - 5% +7Y = -25 — (1) —-@) Multiply equation (I) with 5 we get => 5x32 + 4x54 = -14x5 15 x + 20y=-70 — Multiply equ